# Pricing Options on Forwards in Function-Valued Affine Stochastic Volatility Models

### v1

@@ 2.1.1. A pure-jump $\mathcal{H}^{+}$-valued instantaneous covariance process
      $\int_{\mathcal{H}^{+}\setminus\{0\}}\|\xi\|_{2}^{2}\,m(\mathop{}\!\textnormal{d}\xi)<\infty$ and
@@
      $\int_{\mathcal{H}^{+}\setminus\{0\}}|\langle\chi(\xi),h\rangle|\,m(\mathop{}\!\textnormal{d}\xi)<\infty$ for all $h\in\mathcal{H}$ and there exists an element $I_{m}\in\mathcal{H}$ such that $\langle I_{m},h\rangle=\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),h\rangle\,m(\mathop{}\!\textnormal{d}\xi)$ for every $h\in\mathcal{H}$ ;
@@
    $\langle b,v\rangle-\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),v\rangle\,m(\mathop{}\!\textnormal{d}\xi)\geq 0\,\quad\text{for all}\;v\in\mathcal{H}^{+}\,;$ (2.1)
@@
    $\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),u\rangle\frac{\langle\mu(\mathop{}\!\textnormal{d}\xi),x\rangle}{\|\xi\|_{2}^{2}}<\infty,$
@@
    $\left\langle B^{*}(u),x\right\rangle-\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),u\rangle\frac{\langle\mu(\mathop{}\!\textnormal{d}\xi),x\rangle}{\|\xi\|_{2}^{2}}\geq 0,$
@@
Assume that $(b,B,m,\mu)$ is an admissible parameter such that $\mu(\mathop{}\!\textnormal{d}\xi)$ takes values in $\mathcal{L}_{1}(H)\cap\mathcal{H}^{+}$. Moreover, assume that either one of the following two cases holds:
@@
    $\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}\leq 1\}}\|\xi\|_{2}\,m(\mathop{}\!\textnormal{d}\xi)<\infty\,$ and $\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}\leq 1\}}\|\xi\|_{2}^{-1}\langle y,\mu(\mathop{}\!\textnormal{d}\xi)\rangle<\infty$, $\forall y\in\mathcal{H}^{+}.$
Given an admissible parameter set $(b,B,m,\mu)$ that satisfies Assumptions Assumption $\mathfrak{A}$ and Assumption $\mathfrak{B}$, we conclude from [15, Theorem 2.8] that there exists a probabilistically and analytically weak solution to
$\mathop{}\!\textnormal{d}\mathsf{Y}_{t}$ $=\Big{(}b+B(\mathsf{Y}_{t})+\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}>1\}}\xi\,M(\mathsf{Y}_{t},\mathop{}\!\textnormal{d}\xi)\Big{)}\mathop{}\!\textnormal{d}t+\mathop{}\!\textnormal{d}J_{t},\qquad t\geq 0\,$, (2.2)
@@
$M(y,\mathop{}\!\textnormal{d}\xi):=m(\mathop{}\!\textnormal{d}\xi)+\frac{\langle\mu(\mathop{}\!\textnormal{d}\xi),y\rangle}{\|\xi\|_{2}^{2}}.$ (2.3)
Notice that $(\mathsf{Y}_{t})_{t\geq 0}$ is a square integrable $\mathcal{H}^{+}$-valued Markovian semimartingale and (2.2) amounts to the semimartingale representation of $(\mathsf{Y}_{t})_{t\geq 0}$, see [16, Proposition 2.6] for the details.
@@ 2.1.2. The affine pure-jump joint stochastic volatility model
The first stochastic volatility model we consider takes as instantaneous volatility process the square root of the pure-jump $\mathcal{H}^{+}$-valued process from (2.2). We denote the forward price dynamics by $(X_{t})_{t\geq 0}$, governed by the HJMM–SPDE (1.1), but formulated in a Hilbert space $H$ as follows. Fix an initial value $(x,\mathsf{y})$ with $\mathsf{y}\in\mathcal{H}^{+}$ and $x\in H$, and model parameters $(b,B,m,\mu,D,\Upsilon,\mathcal{A})$. We consider the coupled SDE
$\begin{cases}\mathop{}\!\textnormal{d}X_{t}=\big{(}\mathcal{A}X_{t}+D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}\mathsf{Y}_{t}^{1/2}\,\mathop{}\!\textnormal{d}W_{t},\\ \mathop{}\!\textnormal{d}\mathsf{Y}_{t}=\Big{(}b+B(\mathsf{Y}_{t})+\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}>1\}}\xi\,M(\mathsf{Y}_{t},\mathop{}\!\textnormal{d}\xi)\Big{)}\mathop{}\!\textnormal{d}t+\mathop{}\!\textnormal{d}J_{t},\quad t\geq 0,\end{cases}$ (2.4)
where $(W_{t})_{t\geq 0}$ and $(J_{t})_{t\geq 0}$ are independent. Here $(b,B,m,\mu)$ is an admissible parameter set in the sense of Assumption Assumption $\mathfrak{A}$, $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ denotes the first-derivative (in the maturity variable) operator on the Hilbert space $H$, $D\in\mathcal{L}_{1}(H)$, and $\Upsilon\in H$; see Section 3 for details.
The existence of a (stochastically) weak solution to (2.4) in $H\times\mathcal{H}$ follows from [16, Proposition 2.6 and Lemma 2.8]. As in [16], and motivated by applications, we include the linear term $D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon$ in the drift of $(X_{t})_{t\geq 0}$. We refer to (2.4) as the pure-jump affine stochastic volatility model.
@@ 2.1.3. Riccati equations associated with the pure-jump affine volatility model
Next we aim to derive the Laplace transform of our joint stochastic volatility model. The Laplace will be computed in terms of a solution to ordinary differential equations known as Riccati equations. We start by introducing the Riccati equations associated with our introduced model.
@@
$\mathcal{O}=\left\{u\in\mathcal{H}\colon\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}\,M(y,\mathop{}\!\textnormal{d}\xi)<\infty,\quad\forall y\in\mathcal{H}^{+}\right\}.$
We note that $\mathcal{O}$ is a convex subset in $\mathcal{H}$ and since $\mathcal{H}^{+}\subseteq\mathcal{O}$, it is non-empty. Moreover, we assume that $M(x,\mathop{}\!\textnormal{d}\xi)$ is such that $\mathcal{O}$ is an open subset of $\mathcal{H}$. Define ${\sf F}\colon\mathcal{O}\to\mathbb{R}$ and ${\sf R}\colon H\times\mathcal{O}\to\mathcal{H}$, respectively as
${\sf F}(u)$ $=\langle\tilde{b},u\rangle-\int_{\mathcal{H}^{+}\setminus\{0\}}\big{(}\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}-1\big{)}m(\mathop{}\!\textnormal{d}\xi),$ (2.5)
${\sf R}(h,u)$ $=\tilde{B}^{*}(u)-\tfrac{1}{2}(D^{1/2}h)^{\otimes 2}-D^{1/2}h\otimes D^{1/2}\Upsilon-\int_{\mathcal{H}^{+}\setminus\{0\}}\big{(}\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}-1)\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}},$ (2.6)
where $\tilde{b}\coloneqq b-\int_{\mathcal{H}^{+}\setminus\{0\}}\chi(\xi)\,m(\mathop{}\!\textnormal{d}\xi)$ and for all $v\in\mathcal{H}^{+}$
$\tilde{B}(v)\coloneqq B(v)-\int_{\mathcal{H}^{+}\setminus\{0\}}\chi(\xi)\frac{\langle v,\mu(\mathop{}\!\textnormal{d}\xi)\rangle}{\|\xi\|_{2}^{2}}.$
@@
$\,\frac{\partial\sf\sf P}{\partial t}(t,u)$ $={\sf F}({\sf q}_{2}(t,u)),$  $\,0<t\leq T,\quad{\sf P}(0,u)=0,$ (2.7a)
$\,{\sf q}_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\left(\int_{0}^{t}{\sf q}_{1}(s,u)\mathop{}\!\textnormal{d}s\right),$  $\,0<t\leq T,\quad{\sf q}_{1}(0,u)=u_{1},$ (2.7b)
@@
$\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}|\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}|M(y,\mathop{}\!\textnormal{d}\xi)=\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle\xi,\Re(u)\rangle}M(y,\mathop{}\!\textnormal{d}\xi)<\infty\,,\quad\forall y\in\mathcal{H}^{+}\,.$
@@
$\,\psi_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\mathop{}\!\textnormal{d}s,\quad 0<t\leq T,\quad\psi_{1}(0,u)=u_{1}\,$, (2.8b)
@@
Let $T\geq 0$, $u=(u_{1},0)$, for $u_{1}\in H^{\mathbb{C}}$. Analogously to the extended generalized Riccati equations (2.7a)-(2.7c), we say that the map $(\phi(\cdot,u),\psi_{1}(\cdot,u),\psi_{2}(\cdot,u))$ is a solution to equations (2.8a)-(2.8c) whenever $\phi(\cdot,u)\in C^{1}([0,T],\mathbb{C}),\psi_{1}(\cdot,u)\in C([0,T],H^{\mathbb{C}}),\psi_{2}(\cdot,u)\in C^{1}([0,T],S(\mathcal{O}))$ satisfy (2.8a)-(2.8c).
@@ 2.1.4. Fourier-Laplace transforms of the pure-jump affine volatility model
From Proposition 5.5 and following similar derivations as in [33, Theorem 2.26 and Section 5.3], we deduce the following result:
@@
Let $(b,B,m,\mu)$ be an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$ and Assumption $\mathfrak{B}$, let $(X,\sf Y)$ be the stochastic volatility model (2.4). Let $u=(u_{1},u_{2})\in H^{\mathbb{C}}\times\mathcal{S}(\mathcal{O})$. Let $(\sf P,\sf q_{1},\sf q_{2})$ be the mild solution to the extended Riccati equations (2.7a)-(2.7c) on $[0,T]$ with initial value $(\Re(u_{1}),\Re(u_{2}))$, the existence of which is guaranteed by Proposition 5.2. Then it holds $\mathbb{E}\left[|\mathop{}\!\textnormal{e}^{\langle X_{t},u_{1}\rangle_{H^{\mathbb{C}}}}|\right]<\infty$, and
$\mathbb{E}\left[\mathop{}\!\textnormal{e}^{\langle X_{T},u_{1}\rangle_{H^{\mathbb{C}}}}\mid\mathcal{F}_{t}\right]=\mathop{}\!\textnormal{e}^{-\phi(t,u_{1},0)+\langle X_{t},\psi_{1}(T-t,u_{1},0)\rangle_{H^{\mathbb{C}}}-\langle Y_{t},\psi_{2}(T-t,u_{1},0)\rangle_{\mathcal{H}^{\mathbb{C}}}},$

### v2

@@ 2.1.1. A pure-jump $\mathcal{H}^{+}$-valued instantaneous covariance process
      $\int_{\mathcal{H}^{+}\setminus\{0\}}\|\xi\|_{2}^{2}\,m(\mathop{}\!\mathrm{d}\xi)<\infty$ and
@@
      $\int_{\mathcal{H}^{+}\setminus\{0\}}|\langle\chi(\xi),h\rangle|\,m(\mathop{}\!\mathrm{d}\xi)<\infty$ for all $h\in\mathcal{H}$ and there exists an element $I_{m}\in\mathcal{H}$ such that $\langle I_{m},h\rangle=\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),h\rangle\,m(\mathop{}\!\mathrm{d}\xi)$ for every $h\in\mathcal{H}$ ;
@@
    $\langle b,v\rangle-\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),v\rangle\,m(\mathop{}\!\mathrm{d}\xi)\geq 0\,\quad\text{for all}\;v\in\mathcal{H}^{+}\,;$ (2.1)
@@
    $\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),u\rangle\frac{\langle\mu(\mathop{}\!\mathrm{d}\xi),x\rangle}{\|\xi\|_{2}^{2}}<\infty,$
@@
    $\left\langle B^{*}(u),x\right\rangle-\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),u\rangle\frac{\langle\mu(\mathop{}\!\mathrm{d}\xi),x\rangle}{\|\xi\|_{2}^{2}}\geq 0,$
@@
Assume that $(b,B,m,\mu)$ is an admissible parameter such that $\mu(\mathop{}\!\mathrm{d}\xi)$ takes values in $\mathcal{L}_{1}(H)\cap\mathcal{H}^{+}$. Moreover, assume that either one of the following two cases holds:
@@
    $\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}\leq 1\}}\|\xi\|_{2}\,m(\mathop{}\!\mathrm{d}\xi)<\infty\,$ and $\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}\leq 1\}}\|\xi\|_{2}^{-1}\langle y,\mu(\mathop{}\!\mathrm{d}\xi)\rangle<\infty$, $\forall y\in\mathcal{H}^{+}.$
Given an admissible parameter set $(b,B,m,\mu)$ that satisfies Assumptions Assumption $\mathfrak{A}$ and Assumption $\mathfrak{B}$, we conclude from [14, Theorem 2.8] that there exists a probabilistically and analytically weak solution to
$\mathop{}\!\mathrm{d}\mathsf{Y}_{t}$ $=\Big(b+B(\mathsf{Y}_{t})+\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}>1\}}\xi\,M(\mathsf{Y}_{t},\mathop{}\!\mathrm{d}\xi)\Big)\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}J_{t},\qquad t\geq 0\,$, (2.2)
@@
$M(y,\mathop{}\!\mathrm{d}\xi):=m(\mathop{}\!\mathrm{d}\xi)+\frac{\langle\mu(\mathop{}\!\mathrm{d}\xi),y\rangle}{\|\xi\|_{2}^{2}}.$ (2.3)
Notice that $(\mathsf{Y}_{t})_{t\geq 0}$ is a square integrable $\mathcal{H}^{+}$-valued Markovian semimartingale and (2.2) amounts to the semimartingale representation of $(\mathsf{Y}_{t})_{t\geq 0}$, see [15, Proposition 2.6] for the details.
@@ 2.1.2. The affine pure-jump joint stochastic volatility model
The first stochastic volatility model we consider takes as instantaneous volatility process the square root of the pure-jump $\mathcal{H}^{+}$-valued process from (2.2). We denote the forward price dynamics by $(X_{t})_{t\geq 0}$, governed by the HJMM–SPDE (1.1), but formulated in a Hilbert space $H$ as follows. Fix an initial value $(x,\mathsf{y})\in H\times\mathcal{H}^{+}$ and consider the coupled SDE
$\begin{cases}\mathop{}\!\mathrm{d}X_{t}=\big(\mathcal{A}X_{t}+D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon\big)\,\mathop{}\!\mathrm{d}t+D^{1/2}\mathsf{Y}_{t}^{1/2}\,\mathop{}\!\mathrm{d}W_{t},\\ \mathop{}\!\mathrm{d}\mathsf{Y}_{t}=\Big(b+B(\mathsf{Y}_{t})+\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}>1\}}\xi\,M(\mathsf{Y}_{t},\mathop{}\!\mathrm{d}\xi)\Big)\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}J_{t},\quad t\geq 0,\end{cases}$ (2.4)
where $(W_{t})_{t\geq 0}$ is an $H$-cylindrical Brownian motion, $(J_{t})_{t\geq 0}$ is as described after equation (2.2), $(W_{t})_{t\geq 0}$ and $(J_{t})_{t\geq 0}$ are independent, $(b,B,m,\mu)$ is an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$ and Assumption $\mathfrak{B}$, $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ is the generator of a strongly continuous semigroup on $H$ (in applications, the differentiation operator on the Filipović space; see Section 3), $D\in\mathcal{S}_{1}^{+}(H)$, and $\Upsilon\in H$. We include the linear term $D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon$ in the drift of $(X_{t})_{t\geq 0}$, as motivated by the no-arbitrage drift condition in forward markets; see Section 3 for details. This addition does not affect well-posedness or the affine property; see, for example, the proof of [15, Lemma 2.8].
The existence of a (stochastically) weak solution to (2.4) in $H\times\mathcal{H}$ follows from [15, Proposition 2.6 and Lemma 2.8]. We refer to (2.4) as the pure-jump affine stochastic volatility model.
Remark 2.1 (Independence assumption).
The independence of the driving noise $(W_{t})_{t\geq 0}$ of the forward price dynamics and the noise $(J_{t})_{t\geq 0}$ (respectively $(B_{t})_{t\geq 0}$ in the Wishart model of Section 2.2) of the volatility process is a modeling limitation: it precludes the leverage effect, i.e., the empirically observed negative correlation between returns and volatility changes. This assumption is standard in the infinite-dimensional affine framework as it preserves the affine structure. Extending the model to incorporate leverage via correlated driving noises has been done for the BNS-model in [7].
@@ 2.1.3. Riccati equations associated with the pure-jump affine volatility model
Next we aim to derive the Laplace transform of our joint stochastic volatility model. The Laplace transform will be computed in terms of a solution to ordinary differential equations known as Riccati equations. We start by introducing the Riccati equations associated with our introduced model.
@@
$\mathcal{O}=\left\{u\in\mathcal{H}\colon\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\mathrm{e}^{-\langle\xi,u\rangle}\,M(y,\mathop{}\!\mathrm{d}\xi)<\infty,\quad\forall y\in\mathcal{H}^{+}\right\}.$
We note that $\mathcal{O}$ is a convex subset in $\mathcal{H}$ and since $\mathcal{H}^{+}\subseteq\mathcal{O}$, it is non-empty. Moreover, we assume that $M(x,\mathop{}\!\mathrm{d}\xi)$ is such that $\mathcal{O}$ is an open subset of $\mathcal{H}$. Define ${\sf F}\colon\mathcal{O}\to\mathbb{R}$ and ${\sf R}\colon H\times\mathcal{O}\to\mathcal{H}$, respectively as
${\sf F}(u)$ $=\langle\tilde{b},u\rangle-\int_{\mathcal{H}^{+}\setminus\{0\}}\big(\mathop{}\!\mathrm{e}^{-\langle\xi,u\rangle}-1\big)m(\mathop{}\!\mathrm{d}\xi),$ (2.5)
${\sf R}(h,u)$ $=\tilde{B}^{*}(u)-\tfrac{1}{2}(D^{1/2}h)^{\otimes 2}-D^{1/2}h\otimes D^{1/2}\Upsilon-\int_{\mathcal{H}^{+}\setminus\{0\}}\big(\mathop{}\!\mathrm{e}^{-\langle\xi,u\rangle}-1\big)\frac{\mu(\mathop{}\!\mathrm{d}\xi)}{\|\xi\|_{2}^{2}},$ (2.6)
where $\tilde{b}\coloneqq b-\int_{\mathcal{H}^{+}\setminus\{0\}}\chi(\xi)\,m(\mathop{}\!\mathrm{d}\xi)$ and for all $v\in\mathcal{H}^{+}$
$\tilde{B}(v)\coloneqq B(v)-\int_{\mathcal{H}^{+}\setminus\{0\}}\chi(\xi)\frac{\langle v,\mu(\mathop{}\!\mathrm{d}\xi)\rangle}{\|\xi\|_{2}^{2}}.$
@@
$\,\frac{\partial{\sf P}}{\partial t}(t,u)$ $={\sf F}({\sf q}_{2}(t,u)),$  $\,0<t\leq T,\quad{\sf P}(0,u)=0,$ (2.7a)
$\,{\sf q}_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\left(\int_{0}^{t}{\sf q}_{1}(s,u)\mathop{}\!\mathrm{d}s\right),$  $\,0<t\leq T,\quad{\sf q}_{1}(0,u)=u_{1},$ (2.7b)
@@
Remark 2.3 (Local well-posedness on $\mathcal{O}$).
Local Lipschitz continuity of the map $\mathsf{R}$ on $\mathcal{O}$ guarantees short-time existence of the Riccati flow. In this paper we use the Riccati equations only on time horizons for which a solution is known to exist. Proposition 5.2 provides local existence up to the maximal lifetime, while global existence is verified explicitly only in the concrete examples treated later.
@@
$\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}|\mathop{}\!\mathrm{e}^{-\langle\xi,u\rangle}|M(y,\mathop{}\!\mathrm{d}\xi)=\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\mathrm{e}^{-\langle\xi,\Re(u)\rangle}M(y,\mathop{}\!\mathrm{d}\xi)<\infty\,,\quad\forall y\in\mathcal{H}^{+}\,.$
@@
$\,\psi_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\mathop{}\!\mathrm{d}s,\quad 0<t\leq T,\quad\psi_{1}(0,u)=u_{1}\,$, (2.8b)
@@
Let $T\geq 0$, $u=(u_{1},0)$, for $u_{1}\in H^{\mathbb{C}}$. Analogously to the extended generalized Riccati equations (2.7a)-(2.7c), we say that the map $(\phi(\cdot,u),\psi_{1}(\cdot,u),\psi_{2}(\cdot,u))$ is a solution to equations (2.8a)-(2.8c) whenever $\phi(\cdot,u)\in C^{1}([0,T],\mathbb{C}),\psi_{1}(\cdot,u)\in C([0,T],H^{\mathbb{C}}),\psi_{2}(\cdot,u)\in C^{1}([0,T],S(\mathcal{O}))$ satisfies (2.8a)-(2.8c).
@@ 2.1.4. Fourier-Laplace transforms of the pure-jump affine volatility model
From Proposition 5.5 and following similar derivations as in [30, Theorem 2.26 and Section 5.3], we deduce the following result:
@@
Let $(b,B,m,\mu)$ be an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$ and Assumption $\mathfrak{B}$, let $(X,\sf Y)$ be the stochastic volatility model (2.4), and let $u_{1}\in H^{\mathbb{C}}$. Let $(\sf P,\sf q_{1},\sf q_{2})$ be the mild solution to the extended Riccati equations (2.7a)-(2.7c) on $[0,T]$ with initial value $(\Re(u_{1}),0)$, the existence of which is guaranteed by Proposition 5.2. Then it holds $\mathbb{E}\left[|\mathop{}\!\mathrm{e}^{\langle X_{t},u_{1}\rangle_{H^{\mathbb{C}}}}|\right]<\infty$, and
$\mathbb{E}\left[\mathop{}\!\mathrm{e}^{\langle X_{T},u_{1}\rangle_{H^{\mathbb{C}}}}\mid\mathcal{F}_{t}\right]=\mathop{}\!\mathrm{e}^{-\phi(T-t,u_{1},0)+\langle X_{t},\psi_{1}(T-t,u_{1},0)\rangle_{H^{\mathbb{C}}}-\langle Y_{t},\psi_{2}(T-t,u_{1},0)\rangle_{\mathcal{H}^{\mathbb{C}}}},$